Shift report 20010806 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Aug 06 2001 - 07:52:52 EDT

  • Next message: Hironori Ito: "New Si and BB dN/dEta at 100 Gev/A !!!"

    ------------------------------------------------------------
    Date: Mon Aug  6 07:52:52 2001, Shift: 00:00-08:00
    Supervisor: S.Zgura, with: T.Lussand
    Log book#7 pages 30 - 33
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 03:15 till end of shift
       03:23 - Started run# 4001
       05:37 - Ended run because T1 was broken
    	 - Total of 701512 events
    
       05:47 - Started run# 4002 with T1 off
             - Was still running when shift ended at 08:00
    
    3. T1 stopped functioning at 05:35, and was tried restarted 
      after a slowly decreasing of the HV.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Aug 06 2001 - 07:53:48 EDT